Skysong Company had the following select transactions.


Apr 1, 2020 Accepted Goodwin Company's 12-month, 16% note in settlement of a $53,000 account receivable. 

July 1, 2020 Loaned $55,000 cash to Thomas Slocombe on a 9-month, 14% note. 

Dec. 31, 2020 Accrued interest on all notes receivable. 

Apr. 1, 2021 Received principal plus interest on the Goodwin note. 

Apr. 1.2021 Thomas Slocombe dishonored its note; Skysong expects it will eventually collect. 


Prepare journal entries to record the transactions. Skysong prepares adjusting entries once a year on December 31.
Answer #1
Apr. 1,2020 Notes Receivable 53000
       Accounts Receivable 53000
July. 1,2020 Notes Receivable 55000
       Cash 55000
Dec 31,2020 Interest Receivable 6360 =53000*16%*9/12
        Interest Revenue 6360
Dec 31,2020 Interest Receivable 3850 =55000*14%*6/12
        Interest Revenue 3850
Apr 1,2021 Cash 61480
        Interest Receivable 6360
        Interest Revenue 2120 =53000*16%*3/12
        Notes Receivable 53000
Apr 1,2021 Accounts Receivable 60775
        Interest Receivable 3850
        Interest Revenue 1925 =55000*14%*3/12
        Notes Receivable 55000